# Modeling Clinical Uncertainty in Radiology Reports: from Explicit Uncertainty Markers to Implicit Reasoning Pathways

### v1

@@ 5.  Conclusion
In this work, we present the first systematic approach to addressing both explicit and implicit uncertainty in radiology reports. We rigorously quantify the degree of explicit uncertainty of findings in CXR reports, through an LLM-based automated framework. While demonstrated on the Lunguage dataset, this framework can be applied to any CXR report corpus, enabling the enrichment of widely used benchmarks such as CheXpert (chexpert) and MIMIC-CXR (mimic_cxr_jpg) (of which Lunguage++ is only a subset) with continuous uncertainty measures.
In parallel, we expose and address implicit uncertainty in radiology reports arising from omitted elements of diagnostic reasoning. We introduce a rule-based expansion framework based on expert-defined diagnostic pathways for 14 common CXR diagnoses, which add characteristic sub-findings that may have been left unstated in the original reports. The diagnostic pathways can easily be reused for other tasks, while the expansion framework is applicable to any dataset that has been structured using the Lunguage framework proposed by lunguage.
@@
For the explicit uncertainty framework, a key limitation lies in our reliance on LLM-based pairwise comparisons to construct the reference ranking of hedging phrases. While this approach offers the advantage of scalability (enabling 8,610 comparisons across four LLMs, far beyond what would be feasible with human raters) it also introduces dependency on model behavior. Moreover, our mapping from skill level $\mu$ to probability $p$ is based on assessments from only two radiologists; incorporating a larger expert pool would better capture the variability in how uncertainty is interpreted, as noted in prior studies (communication_doubt_radiology; diagnostic_uncertainty_radiology). Finally, the current framework incurs nontrivial costs when assigning probabilities to new sentences, due to repeated LLM comparisons with the reference ranking. Future work could mitigate this by training a lightweight, locally hosted model, fine-tuned on our high-quality LLM comparison data.
@@
This research made use of and expanded the Lunguage dataset (lunguage). This dataset is derived from MIMIC-CXR (mimic_cxr), a public dataset of chest radiographs and free-text radiology reports. As this dataset contains sensitive patient information, we will follow Physionet’s guidelines by publishing Lunguage++ under the same agreement as the source data. Furthermore, to avoid passing sensitive patient data to public LLM APIs, we followed Physionet’s recommendations by (i) running a HIPAA-compliant GPT-4o model provided by Azure, (ii) revoking data retention rights for Gemini and Claude, and (iii) running MedGemma locally.
@@ A.2.  Reference ranking of hedging phrases
When we use an LLM to compare two hedging phrases, we are essentially playing a match where one hedging phrase is the winner. After each match, TrueSkill updates both $\mu$ and $\sigma$ using Bayesian inference: if an item performs better than expected, its $\mu$ increases, with the magnitude of this change depending on $\sigma$, and vice versa if it performs worse than expected. This mechanism is formalized in TrueSkill’s update equations (trueskill), which ensure that $\sigma$ naturally shrinks as the system becomes more confident in its estimate, stabilizing each item’s rating. Since the winner of a match is picked as the phrase that is closest to “finding is certainly present”, a higher $\mu$ will be obtained for phrases closer to that part of the likelihood spectrum. In our case, we input all pairwise comparisons (across 861 pairs of hedging phrases, each repeated 10 times, with 4 LLMs) one after the other, in a random order. Since the final ranking is order-dependent, we repeat this 10 times with different random orders, averaging the final $\mu$ and $\sigma$ across runs.
@@ A.3.  Fitting each finding-sentence pair into the reference ranking
Using Algorithm A1, we can fit any tentative finding-sentence pair into the reference ranking. Here, the draw probability between the target sentence $t_{tar}$ with $(\mu_{tar},\sigma_{tar})$, and an opponent phrase oppopp with $(\mu_{opp},\sigma_{opp})$ is calculated as follows: $DrawProb(t_{tar},opp)=exp({\frac{-(\mu_{tar}-\mu_{opp})^{2}}{2c^{2}}})\sqrt{d}$, with $c^{2}=2\beta^{2}+\sigma_{tar}^{2}+\sigma_{opp}^{2}$, $d=2\frac{\beta^{2}}{c^{2}}$, and $\beta^{2}=25/6$ (trueskill).
@@
To validate our algorithm and the use of draw probability as an opponent selection strategy, we designed the following experiment. While the true rank of individual sentences in our dataset is unknown, we can assess performance using the hedging phrases in our reference vocabulary, whose ranks are known. Specifically, we simulate a modified reference ranking in which one of the 42 hedging phrases is temporarily excluded, then fit that phrase back into the ranking using a modified version of Algorithm A1. We consider two opponent selection strategies: (i) the draw probability strategy described previously, and (ii) a random strategy, where an opponent is chosen uniformly at random at each step. We also test multiple LLM configurations: one using all LLMs for the first $K$ iterations (as in Algorithm A1) and then picking an LLM randomly, and another using a single LLM (Gemini, GPT-4o, Claude, or MedGemma) for all comparisons. Each configuration is repeated with 10 random seeds, which control sentence sampling, LLM selection, and opponent choice in the random setting. At each iteration, we compute the absolute distance between the current estimated rank of the phrase and its true rank in the reference ranking. If the algorithm converges before 100 iterations, the final rank is used for all subsequent iterations. We then average the absolute rank distances across all 43 phrases and 10 seeds. In these experiments, K was set to 10 and N to 5.
@@ B.1.  Diagnostic Pathways
Diagnosis Specific diagnosis Diagnostic pathways
Pleural Effusion Non-loculated pleural effusion; Loculated pleural effusion Non-loculated (one of nine pathways): view: ap, pa, lateral > ent: blunting > status: dp > loc: costophrenic angle && ent: opacity > status: dp > loc: pleural space > attr: hazy, diffuse
Loculated: view: ap, pa, lateral > ent: opacity > status: dp > loc: pleural space > attr: loculated
Pneumothorax Pneumothorax; Tension pneumothorax Pneumothorax: view: ap, pa, lateral > ent: lucency > status: dp > loc: pleural space && ent: marking > status: dn > loc: pulmonary && ent: air > status: dp > loc: lung periphery
Tension: Pneumothorax && ent: shift > status: dp > loc: mediastinal && attr: large amount
Consolidation Consolidation view: ap, pa, lateral > ent: opacity > status: dp > loc: airspace && ent: volume loss > status: dn
Atelectasis Atelectasis view: ap, pa, lateral > ent: opacity > status: dp && ent: volume loss > status: dp
Pneumonia Pneumonia; Aspiration pneumonia; Lobar pneumonia Pneumonia: view: ap, pa, lateral > ent: opacity > status: dp && ent: fever > status: dp
Aspiration: Pneumonia > loc: lung > attr: dependent portion
Lobar: view: ap, pa, lateral > ent: consolidation > status: dp && ent: fever > status: dp
Pulmonary Edema Interstitial pulmonary edema; Alveolar pulmonary edema Interstitial: view: ap, pa, lateral > ent: opacity > status: dp > loc: interstitial
Alveolar: view: ap, pa, lateral > ent: consolidation > status: dp && ent: pleural effusion > status: dp
Bronchitis Bronchitis; Chronic Bronchitis Bronchitis: view: ap, pa, lateral > ent: opacity > status: dp > loc: peribronchial
Chronic: view: ap, pa, lateral > ent: thickening > status: dp > loc: bronchial wall && ent: cough > status: dp
Cardiomegaly Cardiomegaly view: ap, pa > ent: heart size > status: dp > loc: cardiothoracic > attr: increased
CHF CHF view: ap, pa, lateral > ent: cardiomegaly > status: dp && ent: pulmonary edema > status: dp && ent: dyspnea > status: dp
Emphysema Emphysema; Severe emphysema Emphysema: view: ap, pa, lateral > ent: lucency > status: dp > loc: lung parenchyma > attr: diffuse
Severe: Emphysema && ent: pulmonary vascularity > status: dp > attr: decreased && ent: destruction > status: dp > loc: lung parenchyma
COPD COPD view: ap, pa, lateral > ent: hyperinflation > status: dp > loc: lungs && ent: emphysema > status: dp && ent: cough > status: dp
Fracture Acute fracture; Chronic/Old fracture; Healed fracture; Spinal/Compression fracture Acute: view: ap, pa, lateral > ent: disruption > status: dp > loc: bone
Chronic/Old: view: ap, pa, lateral > ent: callus formation > status: dp > loc: bone
Healed: view: ap, pa, lateral > ent: deformity > status: dp > loc: bone
Spinal/Compression: view: ap, pa, lateral > ent: opacity > status: dp > loc: vertebral body > attr: increased && ent: loss of height > status: dp
Tuberculosis Active tuberculosis; Chronic/Non-active tuberculosis Active: view: ap, pa, lateral > ent: opacity > status: dp
Chronic/Non-active: view: ap, pa, lateral > ent: nodules > status: dp > loc: bilateral upper lung > attr: calcified && ent: architectural distortion > status: dp > loc: bilateral upper lung
Lung Cancer Lung Cancer view: ap, pa, lateral > ent: opacity > status: dp > attr: nodular
@@ B.2.1.  Blacklist for Finding Deduplication
We merge near-duplicate findings into a single canonical form using cosine similarity from a clinical embedding model (e.g., BioLord (remy2023biolord)) with a threshold of $0.9$. Even with this high threshold, purely lexical similarity can still spuriously merge clinically distinct statements that differ along critical axes (e.g., left lower lobe new consolidation vs. right lower lobe new consolidation). To avoid false merges of distinct concepts, we apply an explicit blacklist with two layers: (i) exact-pair exclusions and (ii) pattern-level rules. We use a manually curated blacklist of mutually incompatible pairs, which is invoked when two candidates belong to the same study and to the same coarse anatomic region but differ along key discriminative axes such as laterality, lobar/segmental location, diagnostic status, or pathophysiologic mechanism. A candidate pair is merged only if it passes the similarity threshold and does not match any blacklist entry; otherwise both findings are preserved. Representative blacklist entries and patterns are summarized in Table B2.

### v2

@@ 5.  Conclusion
In this work, we present the first systematic approach to addressing both explicit and implicit uncertainty in radiology reports. We rigorously quantify the degree of explicit uncertainty of findings in CXR reports, through an LLM-based automated framework. While demonstrated on the Lunguage dataset, this framework can be applied to any CXR report corpus, enabling the enrichment of widely used benchmarks such as CheXpert (Irvin et al., 2019) and MIMIC-CXR (Johnson et al., 2019b) (of which Lunguage++ is only a subset) with continuous uncertainty measures.
In parallel, we expose and address implicit uncertainty in radiology reports arising from omitted elements of diagnostic reasoning. We introduce a rule-based expansion framework based on expert-defined diagnostic pathways for 14 common CXR diagnoses, which add characteristic sub-findings that may have been left unstated in the original reports. The diagnostic pathways can easily be reused for other tasks, while the expansion framework is applicable to any dataset that has been structured using the Lunguage framework proposed by Moon et al. (2026).
@@
For the explicit uncertainty framework, a key limitation lies in our reliance on LLM-based pairwise comparisons to construct the reference ranking of hedging phrases. While this approach offers the advantage of scalability (enabling 8,610 comparisons across four LLMs, far beyond what would be feasible with human raters) it also introduces dependency on model behavior. Moreover, our mapping from skill level $\mu$ to probability $p$ is based on assessments from only two radiologists; incorporating a larger expert pool would better capture the variability in how uncertainty is interpreted, as noted in prior studies (Hobby et al., 2000; Shinagare et al., 2019). Finally, the current framework incurs nontrivial costs when assigning probabilities to new sentences, due to repeated LLM comparisons with the reference ranking. Future work could mitigate this by training a lightweight, locally hosted model, fine-tuned on our high-quality LLM comparison data.
@@
Data and Code Availability
Lunguage++ and related resources will be made available via Physionet. Code can be found at our Github repository: github.com/prabaey/lunguage_uncertainty.
@@
This research made use of and expanded the Lunguage dataset (Moon et al., 2026). This dataset is derived from MIMIC-CXR (Johnson et al., 2019a), a public dataset of chest radiographs and free-text radiology reports. As this dataset contains sensitive patient information, we will follow Physionet’s guidelines by publishing Lunguage++ under the same agreement as the source data. Furthermore, to avoid passing sensitive patient data to public LLM APIs, we followed Physionet’s recommendations by (i) running a HIPAA-compliant GPT-4o model provided by Azure, (ii) revoking data retention rights for Gemini and Claude, and (iii) running MedGemma locally.
@@ Acknowledgments
Paloma Rabaey’s research is funded by the Research Foundation Flanders (FWO Vlaanderen) with grant number 1170124N. Additional FWO funding was provided in the form of a travel grant, with grant number V414025N. This research also received funding from the Flemish government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme.
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (No.RS-2019-II190075, No.RS-2025-02304967), the Korea Health Industry Development Institute (KHIDI) grant (No.RS-2025-02213750), and National Research Foundation of Korea (NRF) grant (NRF-2020H1D3A2A03100945), funded by the Korean government (MSIT, MOHW).
@@
  • Anthropic (2025) Claude sonnet. Note: https://www.anthropic.com/transparency/model-report Cited by: §3.2.
  • L. Berlin (2000) Pitfalls of the vague radiology report. American Journal of Roentgenology 174 (6), pp. 1511–1518. Cited by: §1, §4.
  • M. A. Bruno, J. Petscavage-Thomas, and H. H. Abujudeh (2017) Communicating uncertainty in the radiology report. American Journal of Roentgenology 209 (5), pp. 1006–1008. Cited by: §1, §1.
  • A. L. Callen, S. M. Dupont, A. Price, B. Laguna, D. McCoy, B. Do, J. Talbott, M. Kohli, and J. Narvid (2020) Between always and never: evaluating uncertainty in radiology reports using natural language processing. Journal of Digital Imaging 33 (5), pp. 1194–1201. Cited by: §1.
  • G. Comanici, E. Bieber, M. Schaekermann, I. Pasupat, N. Sachdeva, I. Dhillon, M. Blistein, O. Ram, D. Zhang, E. Rosen, et al. (2025) Gemini 2.5: pushing the frontier with advanced reasoning, multimodality, long context, and next generation agentic capabilities. arXiv preprint arXiv:2507.06261. Cited by: §3.1, §3.2.
  • J. Gärtner, P. O. Berberat, M. Kadmon, and S. Harendza (2020) Implicit expression of uncertainty–suggestion of an empirically derived framework. BMC Medical Education 20 (1), pp. 83. Cited by: §1, §4.
  • L. R. Goodman (2014) Felson’s principles of chest roentgenology, a programmed text. Elsevier Health Sciences. Cited by: §4.1.
  • R. B. Gunderman (2009) Biases in radiologic reasoning. American Journal of Roentgenology 192 (3), pp. 561–564. Cited by: §1, §4.
  • R. Herbrich, T. Minka, and T. Graepel (2006) TrueSkill™: a bayesian skill rating system. Advances in neural information processing systems 19. Cited by: §A.2, §A.3, §3.2, §3.3.
  • J. L. Hobby, B. Tom, C. Todd, P. Bearcroft, and A. K. Dixon (2000) Communication of doubt and certainty in radiological reports.. The British Journal of Radiology 73 (873), pp. 999–1001. Cited by: §1, §1, §3, Limitations.
  • A. Hurst, A. Lerer, A. P. Goucher, A. Perelman, A. Ramesh, A. Clark, A. Ostrow, A. Welihinda, A. Hayes, A. Radford, et al. (2024) Gpt-4o system card. arXiv preprint arXiv:2410.21276. Cited by: §3.2.
  • J. Irvin, P. Rajpurkar, M. Ko, Y. Yu, S. Ciurea-Ilcus, C. Chute, H. Marklund, B. Haghgoo, R. Ball, K. Shpanskaya, et al. (2019) Chexpert: a large chest radiograph dataset with uncertainty labels and expert comparison. In Proceedings of the AAAI conference on artificial intelligence, Vol. 33, pp. 590–597. Cited by: §1, §1, §1, §5.
  • S. Jain, A. Agrawal, A. Saporta, S. Q. Truong, D. N. Duong, T. Bui, P. Chambon, Y. Zhang, M. P. Lungren, A. Y. Ng, et al. (2021) Radgraph: extracting clinical entities and relations from radiology reports. arXiv preprint arXiv:2106.14463. Cited by: §1.
  • A. E. Johnson, T. J. Pollard, S. J. Berkowitz, N. R. Greenbaum, M. P. Lungren, C. Deng, R. G. Mark, and S. Horng (2019a) MIMIC-cxr, a de-identified publicly available database of chest radiographs with free-text reports. Scientific data 6 (1), pp. 317. Cited by: §2, Ethics Statement.
  • A. Johnson, M. Lungren, Y. Peng, Z. Lu, R. Mark, S. Berkowitz, and S. Horng (2019b) Mimic-cxr-jpg-chest radiographs with structured labels. PhysioNet 101, pp. 215–220. Cited by: §1, §5.
  • S. Khanna, A. Dejl, K. Yoon, S. Q. Truong, H. Duong, A. Saenz, and P. Rajpurkar (2023) Radgraph2: modeling disease progression in radiology reports via hierarchical information extraction. In Machine learning for healthcare conference, pp. 381–402. Cited by: §1.
  • C. S. Lee, P. G. Nagy, S. J. Weaver, and D. E. Newman-Toker (2013) Cognitive and system factors contributing to diagnostic errors in radiology. American Journal of Roentgenology 201 (3), pp. 611–617. Cited by: §1, §4.
  • J. H. Moon, G. Choi, P. Rabaey, M. G. Kim, H. G. Hong, J. O. Lee, H. Yoon, E. Doe, J. Kim, H. Sharma, D. Coelho de Castro, J. Alvarez Valle, and E. Choi (2026) Lunguage: A Benchmark for Structured and Sequential Chest X-ray Interpretation. PhysioNet. Note: Version 1.0.0 External Links: Document, Link Cited by: 3rd item, §1, §1, §2, §4.2, §4.2, §5, Ethics Statement.
  • J. Pearl (2009) Causality. Cambridge university press. Cited by: §4.1.
  • B. I. Reiner (2018) Quantitative analysis of uncertainty in medical reporting: creating a standardized and objective methodology. Journal of digital imaging 31 (2), pp. 145–149. Cited by: §1.
  • F. Remy, K. Demuynck, and T. Demeester (2023) BioLORD-2023: semantic textual representations fusing llm and clinical knowledge graph insights. arXiv preprint arXiv:2311.16075. Cited by: §B.2.1, §4.2.
  • A. Sellergren, S. Kazemzadeh, T. Jaroensri, A. Kiraly, M. Traverse, T. Kohlberger, S. Xu, F. Jamil, C. Hughes, C. Lau, et al. (2025) Medgemma technical report. arXiv preprint arXiv:2507.05201. Cited by: §3.2.
  • A. B. Shinagare, R. Lacson, G. W. Boland, A. Wang, S. G. Silverman, W. W. Mayo-Smith, and R. Khorasani (2019) Radiologist preferences, agreement, and variability in phrases used to convey diagnostic certainty in radiology reports. Journal of the American College of Radiology 16 (4), pp. 458–464. Cited by: §1, §3, Limitations.
  • M. Turner, J. Ive, and S. Velupillai (2021) Linguistic uncertainty in clinical nlp: a taxonomy, dataset and approach. In International Conference of the Cross-Language Evaluation Forum for European Languages, pp. 129–141. Cited by: §1, §1.
  • W. R. Webb and C. B. Higgins (2011) Thoracic imaging: pulmonary and cardiovascular radiology. Lippincott Williams & Wilkins. Cited by: §4.1.
  • J. T. Wu, N. N. Agu, I. Lourentzou, A. Sharma, J. A. Paguio, J. S. Yao, E. C. Dee, W. Mitchell, S. Kashyap, A. Giovannini, et al. (2021) Chest imagenome dataset for clinical reasoning. arXiv preprint arXiv:2108.00316. Cited by: §1.
  • M. Zhang, X. Hu, L. Gu, T. Harada, K. Kobayashi, R. Summers, and Y. Zhu (2023) Cad-chest: comprehensive annotation of diseases based on mimic-cxr radiology report. PhysioNet. Cited by: §1.
@@ A.2.  Reference ranking of hedging phrases
When we use an LLM to compare two hedging phrases, we are essentially playing a match where one hedging phrase is the winner. After each match, TrueSkill updates both $\mu$ and $\sigma$ using Bayesian inference: if an item performs better than expected, its $\mu$ increases, with the magnitude of this change depending on $\sigma$, and vice versa if it performs worse than expected. This mechanism is formalized in TrueSkill’s update equations (Herbrich et al., 2006), which ensure that $\sigma$ naturally shrinks as the system becomes more confident in its estimate, stabilizing each item’s rating. Since the winner of a match is picked as the phrase that is closest to “finding is certainly present”, a higher $\mu$ will be obtained for phrases closer to that part of the likelihood spectrum. In our case, we input all pairwise comparisons (across 861 pairs of hedging phrases, each repeated 10 times, with 4 LLMs) one after the other, in a random order. Since the final ranking is order-dependent, we repeat this 10 times with different random orders, averaging the final $\mu$ and $\sigma$ across runs.
@@ A.3.  Fitting each finding-sentence pair into the reference ranking
Using Algorithm A1, we can fit any tentative finding-sentence pair into the reference ranking. Here, the draw probability between the target sentence $t_{tar}$ with $(\mu_{tar},\sigma_{tar})$, and an opponent phrase oppopp with $(\mu_{opp},\sigma_{opp})$ is calculated as follows: $DrawProb(t_{tar},opp)=exp({\frac{-(\mu_{tar}-\mu_{opp})^{2}}{2c^{2}}})\sqrt{d}$, with $c^{2}=2\beta^{2}+\sigma_{tar}^{2}+\sigma_{opp}^{2}$, $d=2\frac{\beta^{2}}{c^{2}}$, and $\beta^{2}=25/6$ (Herbrich et al., 2006).
@@
To validate our algorithm and the use of draw probability as an opponent selection strategy, we designed the following experiment. While the true rank of individual sentences in our dataset is unknown, we can assess performance using the hedging phrases in our reference vocabulary, whose ranks are known. Specifically, we simulate a modified reference ranking in which one of the 42 hedging phrases is temporarily excluded, then fit that phrase back into the ranking using a modified version of Algorithm A1. We consider two opponent selection strategies: (i) the draw probability strategy described previously, and (ii) a random strategy, where an opponent is chosen uniformly at random at each step. We also test multiple LLM configurations: one using all LLMs for the first $K$ iterations (as in Algorithm A1) and then picking an LLM randomly, and another using a single LLM (Gemini, GPT-4o, Claude, or MedGemma) for all comparisons. Each configuration is repeated with 10 random seeds, which control sentence sampling, LLM selection, and opponent choice in the random setting. At each iteration, we compute the absolute distance between the current estimated rank of the phrase and its true rank in the reference ranking. If the algorithm converges before 100 iterations, the final rank is used for all subsequent iterations. We then average the absolute rank distances across all 42 phrases and 10 seeds. In these experiments, K was set to 10 and N to 5.
@@ B.2.1.  Blacklist for Finding Deduplication
We merge near-duplicate findings into a single canonical form using cosine similarity from a clinical embedding model (e.g., BioLord (Remy et al., 2023)) with a threshold of $0.9$. Even with this high threshold, purely lexical similarity can still spuriously merge clinically distinct statements that differ along critical axes (e.g., left lower lobe new consolidation vs. right lower lobe new consolidation). To avoid false merges of distinct concepts, we apply an explicit blacklist with two layers: (i) exact-pair exclusions and (ii) pattern-level rules. We use a manually curated blacklist of mutually incompatible pairs, which is invoked when two candidates belong to the same study and to the same coarse anatomic region but differ along key discriminative axes such as laterality, lobar/segmental location, diagnostic status, or pathophysiologic mechanism. A candidate pair is merged only if it passes the similarity threshold and does not match any blacklist entry; otherwise both findings are preserved. Representative blacklist entries and patterns are summarized in Table B1.
@@ B.2.3.  Status Conflict Resolution
Diagnosis Specific diagnosis Diagnostic pathways
Pleural Effusion Non-loculated pleural effusion (one of nine pathways) view: ap, pa, lateral > ent: blunting > status: dp > loc: costophrenic angle && ent: opacity > status: dp > loc: pleural space > attr: hazy, diffuse
Loculated pleural effusion view: ap, pa, lateral > ent: opacity > status: dp > loc: pleural space > attr: loculated
Pneumothorax Pneumothorax view: ap, pa, lateral > ent: lucency > status: dp > loc: pleural space && ent: marking > status: dn > loc: pulmonary && ent: air > status: dp > loc: lung periphery
Tension pneumothorax Pneumothorax && ent: shift > status: dp > loc: mediastinal && attr: large amount
Consolidation view: ap, pa, lateral > ent: opacity > status: dp > loc: airspace && ent: volume loss > status: dn
Atelectasis view: ap, pa, lateral > ent: opacity > status: dp && ent: volume loss > status: dp
Pneumonia Pneumonia view: ap, pa, lateral > ent: opacity > status: dp && ent: fever > status: dp
Aspiration pneumonia Pneumonia > loc: lung > attr: dependent portion
Lobar pneumonia view: ap, pa, lateral > ent: consolidation > status: dp && ent: fever > status: dp
Pulmonary Edema Interstitial pulmonary edema view: ap, pa, lateral > ent: opacity > status: dp > loc: interstitial
Alveolar pulmonary edema view: ap, pa, lateral > ent: consolidation > status: dp && ent: pleural effusion > status: dp
Bronchitis Bronchitis view: ap, pa, lateral > ent: opacity > status: dp > loc: peribronchial
Chronic Bronchitis view: ap, pa, lateral > ent: thickening > status: dp > loc: bronchial wall && ent: cough > status: dp
Cardiomegaly view: ap, pa > ent: heart size > status: dp > loc: cardiothoracic > attr: increased
CHF view: ap, pa, lateral > ent: cardiomegaly > status: dp && ent: pulmonary edema > status: dp && ent: dyspnea > status: dp
Emphysema Emphysema view: ap, pa, lateral > ent: lucency > status: dp > loc: lung parenchyma > attr: diffuse
Severe emphysema Emphysema && ent: pulmonary vascularity > status: dp > attr: decreased && ent: destruction > status: dp > loc: lung parenchyma
COPD view: ap, pa, lateral > ent: hyperinflation > status: dp > loc: lungs && ent: emphysema > status: dp && ent: cough > status: dp
Fracture Acute fracture view: ap, pa, lateral > ent: disruption > status: dp > loc: bone
Chronic/Old fracture view: ap, pa, lateral > ent: callus formation > status: dp > loc: bone
Healed fracture view: ap, pa, lateral > ent: deformity > status: dp > loc: bone
Spinal/Compression fracture view: ap, pa, lateral > ent: opacity > status: dp > loc: vertebral body > attr: increased && ent: loss of height > status: dp
Tuberculosis Active tuberculosis view: ap, pa, lateral > ent: opacity > status: dp
Chronic/Non-active tuberculosis view: ap, pa, lateral > ent: nodules > status: dp > loc: bilateral upper lung > attr: calcified && ent: architectural distortion > status: dp > loc: bilateral upper lung
Lung Cancer view: ap, pa, lateral > ent: opacity > status: dp > attr: nodular